OpenStudy (anonymous):

anna has a fish tank in the shape of a rectangular solid and she plans to put 8 fish in the tank. if she requires 6480 cubic inches of the space and the dimensions of the base of the tank are 24 in by 72 in what should be the height of the fish tank ????

OpenStudy (anonymous):

\[24\times 72\times x =6480\] \[1728\times x =6480\] \[x=\frac{6480}{1728}=3.75\]

OpenStudy (anonymous):

help please - find the length of the longest diagonal of a box that is 3 by 6 by 6

OpenStudy (anonymous):

\[\sqrt{x^2+y^2+z^2}\]

OpenStudy (anonymous):

in this case \[\sqrt{6^2+6^2+8^2}=\sqrt{36+36+64}=\sqrt{136}\]\]

OpenStudy (anonymous):

simplest radical form is \[2\sqrt{34}\] if you need that